Suppose that each firm in a competitive industry has the following costs:

Total Cost: TC= 50+1/2 q^2
Marginal Cost: MC= q

where qq is an individual firm's quantity produced.

The market demand curve for this product is

Demand

QD=160−4PQD=160−4P

where PP is the price and QQ is the total quantity of the good.

Each firm's fixed cost is $_____

What is each firm's variable cost?

q

50+1/2 q

1/2q

1/2q^2

Which of the following represents the equation for each firm's average total cost?

50/q

50+1/2q

50/q+1/2q

1/2q

Complete the following table by computing the marginal cost and average total cost for qq from 5 to 15.

The average total cost is at its minimum when the quantity each firm produces (qq) equals _______

.

Which of the following represents the equation for each firm's supply curve in the short run?

q

50−q

1/2q^2

120−1/2q^2

In the long run, the firm will remain in the market and produce if _______.

Currently, there are 16 firms in the market.

In the short run, in which the number of firms is fixed, the equilibrium price is______

and the total quantity produced in the market is_____units. Each firm produces____

units. (Hint: Total supply in the market equals the number of firms times the quantity supplied by each firm.)

In this equilibrium, each firm makes a profit of $______

. (Note: Enter a negative number if the firm is incurring a loss.)

Firms have an incentive to _____ the market.

In the long run, with free entry and exit, the equilibrium price is _____

, and the total quantity produced in the market is_____units. There arefirms in the market, with each firm producing____

units.
Answer #1

Given the total cost function as TC= 50+1/2 q^2,

1. Each firm's fixed cost is equal to $50

2. Each firm's variable cost is represented by 1/2 q^2. [Option 4 is correct]

3. We can find Average total cost or ATC by dividing TC by q

=> ATC = TC/q = [50+1/2 q^2]/q = 50/q + 1/2 q. Option 3 is correct.

4.

q MC=q ATC=50/q + 1/2 q
5 5 12.50
6 6 11.33
7 7 10.64
8 8 10.25
9 9 10.05
10 10 10.00
11 11 10.04
12 12 10.17
13 13 10.34
14 14 10.57
15 15 10.83

Thus the ATC is minimum at 10.00 when the quantity is equal to 10 units.